### v2

@@ Appendix I Compared with AnovL
Figure 22: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the Fibrous category in the DTD dataset. All results are generated by PILOT.
@@
Figure 23: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the Woven category in the DTD dataset. All results are generated by PILOT.
@@
Figure 24: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the Matted category in the DTD dataset. All results are generated by PILOT.
@@
Figure 25: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the Perforated category in the DTD dataset. All results are generated by PILOT.

### v3

@@ Appendix I Compared with AnovL
Figure 22: Visualization of input images (top), ground‐truth masks (middle), and model‐generated anomaly maps (bottom) for the Fibrous category in the DTD dataset. All results are generated by PILOT.
@@
Figure 23: Visualization of input images (top), ground‐truth masks (middle), and model‐generated anomaly maps (bottom) for the Woven category in the DTD dataset. All results are generated by PILOT.
@@
Figure 24: Visualization of input images (top), ground‐truth masks (middle), and model‐generated anomaly maps (bottom) for the Matted category in the DTD dataset. All results are generated by PILOT.
@@
Figure 25: Visualization of input images (top), ground‐truth masks (middle), and model‐generated anomaly maps (bottom) for the Perforated category in the DTD dataset. All results are generated by PILOT.